Logiweb diagnose of pogave

Up Help

Error in proof of Prop 3.2n_{1} Unification failed Prop 3.2c ⊳ \underline{a} · 0 = 0 ⊳ 0 · \underline{a} = 0 ≫ \underline{a} · 0 = 0 · \underline{a}

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-21.UTC:11:42:14.848103 = MJD-53907.TAI:11:42:47.848103 = LGT-4657606967848103e-6